OpenStudy (anonymous):

given that sin θ= 2/7, calculate tan θ

OpenStudy (anonymous):

mm how is that possible first trying to solve for sin and then calculate tan?

OpenStudy (anonymous):

the answers given are a) 2/(square root of) 53 b) 2/3(square root of) 5 c) 3(sqaure root of) 5 / 2 or d) (square root of) 53 / 2 but if sin = o / h than a = 45 so then tan = 2 / (sqaure root of) 45?

OpenStudy (anonymous):

but that's not one of the answers...

OpenStudy (anonymous):

I'm sorry but I don't understand what are you trying to say.What is this geometry?

OpenStudy (anonymous):

yeah it is. sorry what im trying to do is find the adjacent of the triangle using the info from sin θ = 2/7 because sin = opposite / hypotenuse, so than the opposite side of the triangle would be 2 and the hypotenuse would be 7 right? so the adjacent of the triangle would be the answer from 7^2 - 2^2 which is 45^2, so the adjacent would be √45 and because tan = opposite / adjacent than it would be 2 / √45 but thats not one of the four answers...

OpenStudy (anonymous):

First, find cos x by using the trig identity: sin^2 x + cos^2 x = 1 Next, find tan x by the relation tan x = sin x/cos x. a) cos^2 x = 1 - sin^2 x = 1 - 4/49 = 45/49 --> cos x = (3V5)/7 b) tan x = sin x/cos x = (2/7)/(3V5)/7 = 2/3V5 = 2V5/15

OpenStudy (anonymous):

omg thank you *-*